# A New Framework for Modelling Liquidity Pools as Mean Field Games

### 17 Mar 2026 v2

@@ 2 Model Formulation
We consider $N$ traders who trade their two tokens ETH and USDC in a liquidity pool. Let $X_{t}^{i}$ and $Y_{t}^{i}$, $i=1\dots,N$, be the inventories of ETH and USDC, respectively, for the $i$-th trader at time $t$. The dynamics of the ETH inventory for the $i$-th trader are defined as
@@
where $\alpha_{t}^{i}:[0,T]\to\mathbb{R}$ represents the trading rate, these will be the controls; $\sigma^{i}$ represent the respective volatilities, which for simplicity we assume are independent of $i$; and $\sigma^{i}dW_{t}^{i}$ represents a random change in the trader’s inventory, associated with exogenous market events or inherent wallet usage, where $W_{t}^{i}$ are $N$ uncorrelated Wiener processes.
@@
According to the dynamics of ETH inventories for $N$ agents with $N\in\mathbb{N}$, we model the balance of ETH in the pool as follows
@@
where $X_{0}$ is the number of tokens provided by the liquidity provider at an initial time. A detail to consider is that in a continuous-time trading model with multiple agents, it is unclear how simultaneous trades should be handled. More realistic are continuous-time discrete trading models. In the latter, it is reasonable to assume that agents never trade tokens simultaneously, given that there is a continuum of times to choose from. In general, considering the average is a common practice in Mean Field Games theory, since when $N$ tends to infinity, the average tends to the empirical distribution of the controls.
@@
  • Finally, let $\mathcal{P}(A)$ be the space of probability functions over $A$ along with the weak topology $\tau(A)$ that makes the map $q\to\int_{A}fdq$ continuous for each $f\in B(A)$.
@@
Intuitively, if $N$ is large, due to the symmetry of the model, the contribution of player $i$ to the empirical distribution of controls $\hat{q}_{t}^{N}$ is negligible, and $\hat{q}_{t}^{N}$ can be treated as fixed. We consider the limit when the number of agents $N$ tends to infinity, which simplifies the analysis by reducing the impact of individual fluctuations and focusing on the mean dynamics. Formally, we study the behavior of
@@
Given flows $\bm{\mu}=(\mu_{t})_{0\leq t\leq T}$ and $\bm{q}=(q_{t})_{0\leq t\leq T}$ of probability measures on $\mathbb{R}$ and on $A$ resp., and given a control $\alpha\in\mathbb{A}$ we define the associated expected reward by
@@ 3 Mean Field Game solution
    The control space $A$ is a compact and convex subset of $\mathbb{R}$ and the drift $b:[0,T]\times\mathbb{R}\times\mathcal{P}_{\psi}(\mathbb{R})\times A\to\mathbb{R}$ is continuous.
@@
    The quantities of tokens traded by the agents cannot be infinite since $A$ is bounded, i.e., there exists $M$ such that $|\alpha(t,\omega)|\leq M$, $\forall(t,\omega)\in[0,T]\times\Omega$. The control space $A$ is a closed interval containing the origin, therefore it is compact and convex. The drift is a continuous function, as it is the identity.
@@
    The function $\sigma^{-1}b(a)=\sigma^{-1}a$ is uniformly bounded since $A$ is compact. Moreover, since $\psi_{0}(x)=e^{p|x|}$, $\lambda_{0}$ is such that $\int_{\mathbb{R}^{d}}\psi_{0}(x)^{2}\lambda_{0}(dx)<\infty$ and $\sigma$ is constant, the hypotheses of Lemma 5.2 of [4] are verified.
For (S.3) and (S.4), we need to ensure the model is well-posed. Given that the set of admissible controls $A$ is compact we can take the bound used in (S.1) as $M=\displaystyle\max_{a\in A}|a|$. Then, we have
@@
respectively. Note that by assumption (S.5), $A(t,x,\mu,q,z)$ does not depend on $q$, so we will use the notation $A(t,x,\mu,z)$. Furthermore, by assumptions (S.1) and (S.3), the compactness of $A$ and the continuity of $h$ in the variable $a$, we have that $A(t,x,\mu,z)$ is non-empty.
@@ 4 Approximation to Nash
In this section, we shift our focus to the finite-player version of the game and study the convergence of the system towards a Nash equilibrium as the number of players increases. Specifically, we aim to show that the solution obtained from the mean field game provides a foundation for constructing approximate Nash equilibria for the original $N$-player game.
@@ 5 Numerical Analysis and Practical Implications
To isolate the fixed–point structure and validate the numerical MFG pipeline, we run throughout this section a tractable aggregate–impact proxy in which the population effect enters through the empirical mean control. This specification preserves the mean field coupling while avoiding additional nonlinear feedback effects induced by the AMM reserve dynamics. Accordingly, the experiments below should be read as algorithmic and game–theoretic sanity checks (stability, $\varepsilon$–Nash, and finite–$N$ convergence) rather than a calibrated quantitative study of AMM price impact. A fully endogenous AMM–feedback (“realistic impact”) implementation is left for future work.
@@
When $\lambda=\eta=\rho=0$, the one–step value difference becomes nearly flat in the control variable, so the best–response map provides no robust selection mechanism. As a result, the fixed point computation is ill-conditioned: the solver converges to an arbitrary admissible profile (typically close to the midpoint of $A$ in symmetric settings), driven mainly by discretization and numerical asymmetries rather than by economic forces. This behavior is consistent with the fact that, in the absence of inventory penalties, terminal incentives, or control costs, the control problem admits a continuum of near-indifferent solutions, hence equilibrium selection is not identifiable without regularization.
@@
To assess whether the mean–field equilibrium $q^{\star}$ provides a good approximation for large but finite populations, we quantify the discrepancy between the empirical average control $\bar{\alpha}^{N}(t)=\frac{1}{N}\sum_{i=1}^{N}\alpha_{i}(t)$ and $q^{\star}$. For a given number of agents $N\in\{25,50,100,150,200\}$, we define the finite–$N$ gap as
@@
Fig. 6 shows that $\|\bar{\alpha}^{N}-q^{\star}\|_{2}$ decays approximately as a power law in $N$, indicating standard propagation–of–chaos scaling.
@@
i.e., the fraction of agent–time controls that hit the boundary of $A$. We observe higher $\mathrm{sat\_rate}$ precisely in configurations where $A$ is tight and/or the cost parameters push $a^{\star}(x)$ outside $A$ (cf. Experiments A and C).
@@
We compare the baseline pool size $X_{0}=10$ with a deeper pool $X_{0}=100$. Because individual trades represent a smaller fraction of the reserves, the price impact of each action is attenuated, leading to a smaller–magnitude equilibria $q^{\star}$ and reduced liquidity depletion. The finite–$N$ gap also decreases faster for the deeper pool, reflecting the lower variance of price shocks relative to market depth.
@@
    Finite–player simulations converge to the mean field solution as the population size increases. The finite–$N$ gap decreases systematically and remains within narrow confidence bands across independent random seeds.
@@ Appendix A Algorithmic Implementation
where $A$ is a finite grid used for discrete optimization.
@@
  • Population size $N$: number of agents used to approximate the mean field.
  • Time discretization: number of steps and $dt$.
  • Action grid $A$: size and resolution of admissible controls.

### 18 Mar 2026 v3

@@ 2 Model Formulation
We consider NN traders who trade their two tokens ETH and USDC in a liquidity pool. Let $X_{t}^{i}$ and $Y_{t}^{i}$, $i=1\dots,N$, be the inventories of ETH and USDC, respectively, for the $i$-th trader at time $t$. The dynamics of the ETH inventory for the $i$-th trader are defined as
@@
where $\alpha_{t}^{i}:[0,T]\to\mathbb{R}$ represents the trading rate, these will be the controls; $\sigma^{i}$ represent the respective volatilities, which for simplicity we assume are independent of $i$; and $\sigma^{i}dW_{t}^{i}$ represents a random change in the trader’s inventory, associated with exogenous market events or inherent wallet usage, where $W_{t}^{i}$ are NN uncorrelated Wiener processes.
@@
According to the dynamics of ETH inventories for NN agents with $N\in\mathbb{N}$, we model the balance of ETH in the pool as follows
@@
where $X_{0}$ is the number of tokens provided by the liquidity provider at an initial time. A detail to consider is that in a continuous-time trading model with multiple agents, it is unclear how simultaneous trades should be handled. More realistic are continuous-time discrete trading models. In the latter, it is reasonable to assume that agents never trade tokens simultaneously, given that there is a continuum of times to choose from. In general, considering the average is a common practice in Mean Field Games theory, since when NN tends to infinity, the average tends to the empirical distribution of the controls.
@@
  • Finally, let $\mathcal{P}(A)$ be the space of probability functions over AA along with the weak topology $\tau(A)$ that makes the map $q\to\int_{A}fdq$ continuous for each $f\in B(A)$.
@@
Intuitively, if NN is large, due to the symmetry of the model, the contribution of player $i$ to the empirical distribution of controls $\hat{q}_{t}^{N}$ is negligible, and $\hat{q}_{t}^{N}$ can be treated as fixed. We consider the limit when the number of agents NN tends to infinity, which simplifies the analysis by reducing the impact of individual fluctuations and focusing on the mean dynamics. Formally, we study the behavior of
@@
Given flows $\bm{\mu}=(\mu_{t})_{0\leq t\leq T}$ and $\bm{q}=(q_{t})_{0\leq t\leq T}$ of probability measures on $\mathbb{R}$ and on AA resp., and given a control $\alpha\in\mathbb{A}$ we define the associated expected reward by
@@ 3 Mean Field Game solution
    The control space AA is a compact and convex subset of $\mathbb{R}$ and the drift $b:[0,T]\times\mathbb{R}\times\mathcal{P}_{\psi}(\mathbb{R})\times A\to\mathbb{R}$ is continuous.
@@
    The quantities of tokens traded by the agents cannot be infinite since AA is bounded, i.e., there exists $M$ such that $|\alpha(t,\omega)|\leq M$, $\forall(t,\omega)\in[0,T]\times\Omega$. The control space AA is a closed interval containing the origin, therefore it is compact and convex. The drift is a continuous function, as it is the identity.
@@
    The function $\sigma^{-1}b(a)=\sigma^{-1}a$ is uniformly bounded since AA is compact. Moreover, since $\psi_{0}(x)=e^{p|x|}$, $\lambda_{0}$ is such that $\int_{\mathbb{R}^{d}}\psi_{0}(x)^{2}\lambda_{0}(dx)<\infty$ and $\sigma$ is constant, the hypotheses of Lemma 5.2 of [4] are verified.
For (S.3) and (S.4), we need to ensure the model is well-posed. Given that the set of admissible controls AA is compact we can take the bound used in (S.1) as $M=\displaystyle\max_{a\in A}|a|$. Then, we have
@@
respectively. Note that by assumption (S.5), $A(t,x,\mu,q,z)$ does not depend on $q$, so we will use the notation $A(t,x,\mu,z)$. Furthermore, by assumptions (S.1) and (S.3), the compactness of AA and the continuity of $h$ in the variable $a$, we have that $A(t,x,\mu,z)$ is non-empty.
@@ 4 Approximation to Nash
In this section, we shift our focus to the finite-player version of the game and study the convergence of the system towards a Nash equilibrium as the number of players increases. Specifically, we aim to show that the solution obtained from the mean field game provides a foundation for constructing approximate Nash equilibria for the original NN-player game.
@@ 5 Numerical Analysis and Practical Implications
To isolate the fixed–point structure and validate the numerical MFG pipeline, we run throughout this section a tractable aggregate–impact proxy in which the population effect enters through the empirical mean control. This specification preserves the mean field coupling while avoiding additional nonlinear feedback effects induced by the AMM reserve dynamics. Accordingly, the experiments below should be read as algorithmic and game–theoretic sanity checks (stability, $\varepsilon$–Nash, and finite–NN convergence) rather than a calibrated quantitative study of AMM price impact. A fully endogenous AMM–feedback (“realistic impact”) implementation is left for future work.
@@
When $\lambda=\eta=\rho=0$, the one–step value difference becomes nearly flat in the control variable, so the best–response map provides no robust selection mechanism. As a result, the fixed point computation is ill-conditioned: the solver converges to an arbitrary admissible profile (typically close to the midpoint of AA in symmetric settings), driven mainly by discretization and numerical asymmetries rather than by economic forces. This behavior is consistent with the fact that, in the absence of inventory penalties, terminal incentives, or control costs, the control problem admits a continuum of near-indifferent solutions, hence equilibrium selection is not identifiable without regularization.
@@
To assess whether the mean–field equilibrium $q^{\star}$ provides a good approximation for large but finite populations, we quantify the discrepancy between the empirical average control $\bar{\alpha}^{N}(t)=\frac{1}{N}\sum_{i=1}^{N}\alpha_{i}(t)$ and $q^{\star}$. For a given number of agents $N\in\{25,50,100,150,200\}$, we define the finite–NN gap as
@@
Fig. 6 shows that $\|\bar{\alpha}^{N}-q^{\star}\|_{2}$ decays approximately as a power law in NN, indicating standard propagation–of–chaos scaling.
@@
i.e., the fraction of agent–time controls that hit the boundary of AA. We observe higher $\mathrm{sat\_rate}$ precisely in configurations where AA is tight and/or the cost parameters push $a^{\star}(x)$ outside AA (cf. Experiments A and C).
@@
We compare the baseline pool size $X_{0}=10$ with a deeper pool $X_{0}=100$. Because individual trades represent a smaller fraction of the reserves, the price impact of each action is attenuated, leading to a smaller–magnitude equilibria $q^{\star}$ and reduced liquidity depletion. The finite–NN gap also decreases faster for the deeper pool, reflecting the lower variance of price shocks relative to market depth.
@@
    Finite–player simulations converge to the mean field solution as the population size increases. The finite–NN gap decreases systematically and remains within narrow confidence bands across independent random seeds.
@@ Appendix A Algorithmic Implementation
where AA is a finite grid used for discrete optimization.
@@
  • Population size NN: number of agents used to approximate the mean field.
  • Time discretization: number of steps and dtdt.
  • Action grid AA: size and resolution of admissible controls.